A 2.5-kg object experiences a single, time-dependent force Fx(t) = (18 N/s )t - (36 N ).

1.What is the object's acceleration at t = 0?

2.What is the object's acceleration at t = 2.0 s?

3.What is the object's acceleration at t = 4.0 s?
Answer #1

here,

mass of the object , m = 2.5 kg

Fx = (18 t - 36) N

accelration of the object , a = F/m

a = ( 7.2 *t - 14.4 ) m/s^2

1.

the object's acceleration at t = 0 , a0 = -14.4 m/s^2

2.

the object's acceleration at t = 2.0 s , a2 = 0 ms/^2

3.

the object's acceleration at t = 4.0 s , a4 = 7.2 * 4 - 14.4

a4 = 14.4 m/s^2

the object's acceleration at t = 4.0 s is 14.4 m/s^2